# Optimal discrimination of quantum sequences

### v1

@@ Abstract
A key concept of quantum information theory is that accessing information encoded in a quantum system requires us to discriminate between several possible states the system could be in. A natural generalization of this problem, namely, quantum sequence discrimination, appears in various quantum information processing tasks, the objective being to determine the state of a finite sequence of quantum states. Since such a sequence is a composite quantum system, the fundamental question is whether an optimal measurement is local, i.e., comprising measurements on the individual members, or collective, i.e. requiring joint measurement(s). In some known instances of this problem, the optimal measurement is local, whereas in others, it is collective. But, so far, a definite prescription based solely on the problem description has been lacking. In this paper, we prove that if the members of a given sequence are drawn secretly and independently from an ensemble or even from different ensembles, the optimum success probability is achievable by fixed local measurements on the individual members of the sequence, and no collective measurement is necessary. This holds for both minimum-error and unambiguous state discrimination paradigms.
@@
One of the characteristic features of quantum theory is that composite quantum systems can possess nonlocal properties. This is often associated with entangled systems violating a Bell inequality bell1966problem ; brunner2014bell . However, an unentangled system, whose parts may not have interacted in the past, can also exhibit nonlocality, conceptually different from the Bell type peres1991optimal ; bennett1999quantum ; wootters2006distinguishing ; chitambar2013revisiting ; halder2019strong . Specifically, when determining the state of an unentangled system, known to be in one of several possible states, a joint measurement of the whole system is sometimes necessary. Thus a fundamental question in quantum information is when does optimal extraction of information from a composite system demands access to the whole and when it does not.
@@
This question frequently arises in many quantum information theoretic protocols of practical importance, such as quantum change point detection sentis2016quantum ; sentis2017exact ; mohan2023generalized , multiple copy state discrimination higgins2011multiple , and quantum key distribution cerutti2024optimal . In all these scenarios, essentially, the objective is to determine the state of an unknown quantum sequence $\left(\rho_{1},\dots,\rho_{k}\right)$ whose members $\rho_{i}$ are drawn from ensembles of which we have complete knowledge. For a given sequence of finite length, this boils down to the problem of sequence discrimination, a natural generalization of the well-studied quantum state discrimination problem bae2015quantum ; barnett2009quantum ; chefles2000quantum ; bergou200411 , where one aims to discriminate between the possible states a quantum system could be in. Since the state of a sequence $\left(\rho_{1},\dots,\rho_{k}\right)$ is of the form $\rho_{1}\otimes\cdots\otimes\rho_{k}$, the question is whether the optimal measurement, optimal according to some well-defined figure of merit, is local, i.e., comprising measurements on the individual members or collective, i.e. requiring joint measurement(s). Moreover, even when local access is sufficient, it is necessary to know whether coordinated or adaptive strategies provide any advantage over fixed strategies that do not involve adaptation based on the outcomes of already performed measurements. In fact, we know of instances where the optimal measurement is local and fixed gupta2024unambiguous , local and adaptive acin2005multiple , and collective calsamiglia2010local ; in particular, in the quantum change-point problem, it is collective sentis2016quantum , whereas in multiple-copy discrimination, depending on the problem specification, it could be either of the two acin2005multiple ; calsamiglia2010local .
@@
Unfortunately, given a sequence discrimination problem, we do not have any definite characterization or prescription that could, at the very least, tell us about the nature of the optimal measurement. As it is, discriminating between states of a composite quantum system is known to be challenging wootters2006distinguishing ; chitambar2013revisiting ; halder2019strong and for sequences more so as they could be of any finite length and also come in different varieties: they could be i.i.d., where the members of a sequence are independent (i.e., selection of an individual member does not depend on what states have already appeared in the earlier positions) and identically distributed (i.e., the members are drawn from the same ensemble), or non-i.i.d., where members are either drawn independently but from different ensembles or not independently at all.
@@
In this paper, we take a significant step towards addressing whether an optimal measurement is local or collective only from the problem description. We prove that if the members of a sequence are selected secretly and independently either from the same ensemble or even from different ensembles, the optimal measurement is local; in particular, it constitutes fixed measurements on the individual elements. The result holds for both minimum-error and unambiguous state discrimination paradigms. For the latter, however, additional assumptions are required for a nontrivial result chefles1998unambiguous ; feng2004unambiguous .
@@
The main implication of our result is the following. We have shown that if the members of a given sequence are selected independently, irrespective of the probability distribution associated with the selection being identical or different, the optimal measurement to determine its state is always local and fixed. Therefore, if the independence condition does not hold, all types of measurements mentioned earlier, are possible. Future works, therefore, only need to address the optimality question in the non-independent scenario.
@@
Suppose we are given an unknown sequence of length $k\in\mathbb{N}$ and wish to determine its state as well as possible. The members of the given sequence are secretly and independently drawn from potentially different but known ensembles. The ensemble from which the $i$-th member is drawn is labelled as $\mathcal{E}^{i}$, where
@@
with $\rho_{j}^{i}$ being density operators on $\mathbb{C}^{d_{i}}$ for $d_{i}\geqslant 2$. Thus, for a sequence of length $k$, we have $k$ such ensembles $\mathcal{E}^{1},\dots,\mathcal{E}^{k}$. This way, we can account for all possibilities: every member is drawn from a different ensemble; some are drawn from the same ensemble, in which case, though the labelling of ensembles is different, they are, in fact, identical; every member is drawn from the same ensemble, so here no indexing of ensembles would be necessary. Note that the last one corresponds to i.i.d. sequences, whereas every other case corresponds to independent but not identically distributed sequences.
Thus, given a sequence of length $k$, we only know that the $i$-th member is drawn from $\mathcal{E}^{i}$ and $\eta_{j}^{i}$ is the prior probability of it being $\rho_{j}^{i}$. With this, the ensemble of all possible sequences of length $k$ is given by
@@
The optimality of a measurement subject to a given set of states depends on the choice of measurement strategy. Here, we consider both minimum-error and unambiguous state discrimination paradigms. The former minimizes the average error and applies to any set of states. The corresponding measure is the success probability, the maximum probability that the unknown state is correctly determined holevo1973statistical ; helstrom1969quantum . The latter ivanovic1987differentiate ; dieks1988overlap ; peres1988differentiate , however, applies to sets of states that satisfy a specific condition, and if this condition is met, this approach correctly determines the unknown state with a nonzero probability; for example, a set of pure states can be unambiguously discriminated if and only if they are linearly independent chefles1998unambiguous . Closed-form solutions for both are known in the two-state case helstrom1969quantum ; jaeger1995optimal and in specific instances where the states satisfy certain symmetry properties sun2001optimum . For generic ensembles, finding solutions is generally hard; however, semidefinite programs exist eldar2003designing ; eldar2003semidefinite , and the optimum success probability can be obtained as the output of such programs.
In what follows, we first state the main result and then go on to prove it for both minimum-error and unambiguous discrimination.
@@
Let $p\left(\mathcal{E}\right)$ denote the optimum probability for minimum-error or unambiguous discrimination between the elements of an ensemble $\mathcal{E}$. Then
$p\left(\mathcal{E}_{k}\right)=\prod_{i=1}^{k}p\left(\mathcal{E}^{i}\right).$
Thus the optimal measurement for discriminating between the elements of $\mathcal{E}_{k}$ is local and comprises of individual measurements on the members of the sequence. In particular, for the $i$-th member the local measurement corresponds to the optimal measurement that discriminates between the elements of the ensemble $\mathcal{E}^{i}$. Note that, for unambiguous discrimination, a nontrivial result, i.e., $p\left(\mathcal{E}_{k}\right)>0$ is obtained provided for each ensemble $\mathcal{E}^{i}$, the corresponding elements can be unambiguously discriminated.
Minimum-error discrimination of quantum sequences.—We need the following two results to prove Theorem 1 for this case. The first one is the Holevo–Yuen–Kennedy–Lax Theorem yuen1975optimum ; watrous2018theory , which provides a necessary and sufficient condition for the optimality of a minimum-error discrimination measurement for a given ensemble (note that the optimal measurement is not unique bae2013structure ). We only state the theorem here, the proof can be found in watrous2018theory .
Theorem 2 (Holevo–Yuen–Kennedy–Lax yuen1975optimum ; watrous2018theory ).
@@
The second result we need is the following lemma, whose proof can be found in the supplemental material (SM).
@@
Let $A,B,C,$ and $D$ be positive semidefinite operators such that $A\succeq C$ and $B\succeq D$. Then $A\otimes B\succeq C\otimes D$.
@@
Without loss of generality, assume that an optimal measurement for minimum-error discrimination between the elements of the ensemble $\mathcal{E}^{i}$ is $M^{i}=\{M_{x}^{i}\}_{x=1}^{\ell_{i}}\text{ where }M^{i}_{x}\in\operatorname{Pos}(\mathbb{C}^{d_{i}})\text{ and }\sum_{x=1}^{\ell_{i}}M_{x}^{i}=\mathbb{1}_{d_{i}}$ for all $x\in[\ell_{i}]$ and $i\in[k]$, where $\operatorname{Pos}(\mathbb{C}^{d_{i}})$ denotes the set of positive semidefinite operators on $\mathbb{C}^{d_{i}}$. It follows that
@@
$=$ $\mathbb{1}_{d_{1}}\otimes\cdots\otimes\mathbb{1}_{d_{k}},$
showing that $\{M^{1}_{x_{1}}\otimes\cdots\otimes M^{k}_{x_{k}}\}$ is a bonafide measurement on $\mathcal{E}_{k}$. This, together with equation (5), demonstrates that the measurement whose elements are ${M^{1}_{x_{1}}\otimes\cdots\otimes M^{k}_{x_{k}}}$ is indeed an optimal measurement for discriminating the elements of the ensemble $\mathcal{E}_{k}$, as established by Theorem 2.
@@
Unambiguous discrimination of quantum sequences.— Unlike the minimum-error case, for non-trivial unambiguous discrimination, a set of quantum states $Q=\{\sigma_{1},\ldots,\sigma_{N}\}$ has to satisfy the following condition.
Let $\text{supp}(Q)$ denote the Hilbert space spanned by the eigenvectors of the matrices $\{\sigma_{1},\ldots,\sigma_{N}\}$ that correspond to nonzero eigenvalues. Additionally, let $S(\mathcal{E})$ represent the set of states of an ensemble $\mathcal{E}$.
@@
(from feng2004unambiguous ) The set of density operators $Q=\{\sigma_{1},\ldots,\sigma_{N}\}$ can be unambiguously discriminated if and only if $\text{supp}(Q)\neq\text{supp}(Q_{i})$ where $Q_{i}=Q\setminus\{\sigma_{i}\}$ for all $i\in[N]$. If $\sigma_{i}$ are rank $1$ operators (pure states), then this condition is same as the set $Q$ being linearly independent.
Since we now want to discriminate the set of sequences unambiguously, they must satisfy Lemma 4. That is, for $S(\mathcal{E}_{k})$ to be unambiguously discriminable it must hold that
@@
for all $x_{i}\in[\ell_{i}]$ and $i\in[k]$. In the following lemma (proof in SM), we show that a set of quantum sequences satisfies this condition if and only if the individual ensembles do.
@@
The following theorem is immediate from Lemmas 4 and 5.
Theorem 6.
We now present the proof of Theorem 1 for the unambiguous case. The unambiguous discrimination problem of mixed quantum states can be cast as a semidefinite program (SDP) eldar2004optimal . Let us first present this SDP for an ensemble of states $\mathcal{E}=(Q,q)$, where $Q=\{\sigma_{i}\}_{i=1}^{N}$ is a set of density operators acting on $\mathbb{C}^{d}$ and $q=(q_{1},\ldots,q_{N})$ is the vector of a priori probabilities of states of $Q$. Without loss of generality, we can assume that the eigenvectors of the states of $Q$ that correspond to nonzero eigenvalues span $\mathbb{C}^{d}$. This is analogous to assuming that for unambiguous discrimination of pure states, we can take the dimension of the Hilbert space they live in to equal the dimension of the space spanned by them. We define $\widetilde{Q_{i}}$ as the intersection of all kernels $\mathcal{K}_{j}$ of $\sigma_{j}$, excluding $\mathcal{K}_{i}$. This can be expressed mathematically as $\widetilde{Q_{i}}=\cap_{j=1,j\neq i}^{N}\mathcal{K}_{j}$. We also introduce $\Theta_{i}$, a $d\times r_{i}$ matrix whose columns form an orthonormal basis for $\widetilde{Q_{i}}$, where $r_{i}$ is the dimension of $\widetilde{Q_{i}}$. With these definitions in place, we can formulate the SDP for determining the optimal probability of unambiguous discrimination of the ensemble $\mathcal{E}$ as follows
Primal problem
maximize: $\sum_{i=1}^{N}q_{i}\operatorname{Tr}(\sigma_{i}\Theta_{i}\Delta_{i}\Theta_{i}^{\dagger})$
subject to: $\sum_{i=1}^{N}\Theta_{i}\Delta_{i}\Theta_{i}^{\dagger}\preceq\mathbb{1},$ (8a)
$\Delta_{i}\succeq 0,\quad\forall i\in[N]$
Dual problem
minimize: $\operatorname{Tr}(Z)$
subject to: $\Theta_{i}^{\dagger}\left(Z-q_{i}\sigma_{i}\right)\Theta_{i}\succeq 0\quad\forall i\in[N],$
$Z\succeq 0,$ (8b)
where $\Delta_{i}$ is an $r_{i}\times r_{i}$ matrix for each $i\in[N]$. The optimization variables in this formulation are the $N$ matrices denoted by $\Delta_{i}$. Our proof strategy will be to use Slater’s theorem which states that if the primal problem is convex and strict feasibility holds, then the duality gap is zero and the primal and dual optimal values are equal boyd2004convex . We’ll first assume primal and dual optimal variables for the ensemble SDPs. Then we will use these variables to construct primal and dual feasible solutions for the sequence SDP. Finally we will show that these variables make the primal and dual value equal. Applying Slater’s theorem to the sequence SDP, we will conclude that these must be the optimal solutions.
Proof.
Consider the ensemble
where $\rho_{j}^{i}$ are $d_{i}\times d_{i}$ density operators and $j\in[\ell_{i}]$. Let $S^{i}_{j}=S(\mathcal{E}^{i})\setminus\{\rho^{i}_{j}\}$ for some $j\in[\ell_{i}]$ and $\widetilde{S^{i}_{j}}$ be the intersection of the kernels of all the density matrices of $\mathcal{E}^{i}$, except for $\rho^{i}_{j}$. That is, $\widetilde{S^{i}_{j}}=\cap_{t=1,t\neq j}^{\ell_{i}}\mathcal{K}_{t}^{i}$ where $\mathcal{K}_{t}^{i}$ is the kernel of $\rho_{t}^{i}$. Let $\Theta_{j}^{i}$ be a $d_{i}\times r_{j}^{i}$ matrix whose columns form an arbitrary orthonormal basis for $\widetilde{S^{i}_{j}}$ (which is of dimension $r_{j}^{i}$). The optimal probability of unambiguous discrimination of the elements of $\mathcal{E}^{i}$ is denoted by $p(\mathcal{E}^{i})$ and let $\Delta_{j}^{i}$ be the $r_{j}^{i}\times r_{j}^{i}$ matrices that achieve this optimum. In addition to being positive, these matrices satisfy $\sum_{j=1}^{\ell_{i}}\Theta_{j}^{i}\Delta_{j}^{i}{\Theta_{j}^{i}}^{\dagger}\preceq\mathbb{1}$. (See SM for the SDPs of the individual ensembles and the sequence.) We denote the optimal dual variable for ensemble $\mathcal{E}^{i}$ by $Z_{i}$, which is also a positive matrix and satisfies ${\Theta_{j}^{i}}^{\dagger}\left(Z_{i}-\eta_{j}^{i}\rho_{j}^{i}\right)\Theta_{j}^{i}\succeq 0$. The optimal primal and dual variables satisfy
Now, we turn to the sequence ensemble given by equation (2) and consider the SDP of its unambiguous discrimination. Note that there are $\ell$ number of sequence states in this ensemble, and a state is labeled by the index $(x_{1},\ldots,x_{k})$, which is a $k$-tuple of indices where $x_{i}\in[\ell_{i}]$ for all $i\in[k]$. Let us denote by $\widetilde{S}(x_{1},\ldots,x_{k})$ the intersection of the kernels of all the states of $\mathcal{E}_{k}$ except for $\rho^{1}_{x_{1}}\otimes\cdots\otimes\rho^{k}_{x_{k}}$. That is,
where $\mathcal{K}(y_{1},\ldots,y_{k})$ is the kernel of $\rho^{1}_{y_{1}}\otimes\cdots\otimes\rho^{k}_{y_{k}}$. This subspace has a decomposition (see Theorem 9 in SM) as follows
We first give a prescription to construct matrices $\Theta(x_{1},...,x_{k})$ whose columns form an orthonormal basis of $\widetilde{S}(x_{1},\ldots,x_{k})$. From equation (Proof.), it can be seen that these matrices are of size $\prod_{i=1}^{k}d_{i}\times\prod_{i=1}^{k}r^{i}_{x_{i}}(x_{i}\in[\ell_{i}])$ and a simple procedure to construct them is to take $\Theta(x_{1},...,x_{k})=\Theta_{x_{1}}^{1}\otimes\cdots\otimes\Theta_{x_{k}}^{k}$. For a feasible primal variable $\Delta(x_{1},...,x_{k})$, we take the tensor product of the optimal primal variables, $\Delta(x_{1},...,x_{k})=\Delta_{x_{1}}^{1}\otimes\cdots\otimes\Delta_{x_{k}}^{k}$. This operator is positive since it is the tensor product of positive operators. Also, the defined operators satisfy (using Lemma 3),
$\preceq$  $\;\mathbb{1}.$
Now consider the dual variable $Z=Z_{1}\otimes\cdots\otimes Z_{k}$, which is positive due to the positivity of the operators $Z_{i}$ for all $i\in[k]$. We also know that these variables satisfy ${\Theta_{x_{i}}^{i}}^{\dagger}Z_{i}\Theta_{x_{i}}^{i}\succeq{\Theta_{x_{i}}^{i}}^{\dagger}(\eta_{x_{i}}^{i}\rho_{x_{i}}^{i})\Theta_{x_{i}}^{i}$ for all $x_{i}\in[\ell_{i}]$ and $i\in[k]$. By taking tensor products for the indices $(x_{1},...,x_{k})$ we get (using Lemma 3)
Conclusions.—Sequence discrimination problems arise in various quantum information processing tasks, where the objective is to determine the state of a sequence of finite length. The main question here is whether an optimal measurement is local or collective, i.e., whether it suffices to measure the individual members of the sequence or not. In this paper, we showed that as long as the members of the sequences are drawn independently from the same ensemble or even from different ensembles, the optimal measurement in both minimum-error and unambiguous discrimination paradigms is local and comprises fixed measurements on the individual components. It follows that if we give up the condition of independent selection of the members of a sequence, the optimal measurement could be either local and fixed, local and adaptiveacin2005multiple , or collective (partially or wholly) calsamiglia2010local .
Our result also shows that independent sequences do not exhibit nonlocality in the sense some other unentangled systems do, as collective measurement is not necessary for optimal extraction of information from such sequences. However, in the context of quantum state exclusion, one may consider them nonlocal, for an entangled measurement is necessary, as demonstrated by the PBR result pusey2012reality and its generalizations crickmore2020unambiguous .
A significant part of our proof relied on solving the SDP formulation of the problem using the strong duality theorem boyd2004convex . While SDP has been successfully used to solve various problems in quantum information theory skrzypczyk2023semidefinite ; mironowicz2024semi , we believe our approach could help to solve problems in many-body unentangled systems.
Our result could be immediately applied to settings of quantum key distribution protocols Bennett_2014 ; bennett1992quantum or similar ones. In these protocols, Alice sends a sequence of quantum states - selected independently from a known ensemble - to Bob, who measures each state individually. However, if Bob can store the incoming states in quantum memory, he may think about performing a collective measurement on the entire sequence to extract more information. Our result rules out this possibility, as these sequences are i.i.d.
A natural direction for future research involves doing away with the independence assumption. Can there be sequences of dependent quantum states whose optimal discrimination is still achievable by fixed, local measurements? A systematic study involving the set of states, the nature of the joint probability distribution for the states of a sequence and the nature of optimal measurement to discriminate them will be interesting to investigate.
Acknowledgements.
Appendix A Supplemental material
Appendix B Proof of Lemma 3
Proof.
$=(A\otimes(B-D))+((A-C)\otimes D).$
Since $A-C$ and $B-D$ are positive semidefinite, their tensor products with any positive semidefinite operator also yield positive semidefinite operators. Hence, $A\otimes(B-D)$ and $(A-C)\otimes D$ are positive semidefinite. The sum of two positive semidefinite operators remains positive semidefinite. Therefore, $A\otimes B-C\otimes D$ is positive semidefinite, thus $A\otimes B\succeq C\otimes D$. ∎
Appendix C Proof of Lemma 5
We will need the following two lemmas to prove our claim.

### v2

@@ Abstract
A key concept of quantum information theory is that accessing information encoded in a quantum system requires us to discriminate between several possible states the system could be in. A natural generalization of this problem, namely, quantum sequence discrimination, appears in various quantum information processing tasks, the objective being to determine the state of a finite sequence of quantum states. Since such a sequence is a composite quantum system, the fundamental question is whether an optimal measurement is local, i.e., comprising measurements on the individual members, or collective, i.e. requiring joint measurement(s). In some known instances of this problem, the optimal measurement is local, whereas in others, it is collective. But, so far, a definite prescription based solely on the problem description has been lacking. In this paper, we prove that if the members of a given sequence are secretly and independently drawn from an ensemble or even from different ensembles, the optimum success probability is achievable by fixed local measurements on the individual members of the sequence, and no collective measurement is necessary. This holds for both minimum-error and unambiguous state discrimination paradigms.
I Introduction
@@
One of the characteristic features of quantum theory is that composite quantum systems can possess nonlocal properties. This is often associated with entangled systems violating a Bell inequality [1, 2]. However, an unentangled system, whose parts may not have interacted in the past, can also exhibit nonlocality, conceptually different from the Bell type [3, 4, 5, 7, 6]. Specifically, when determining the state of an unentangled system, known to be in one of several possible states, a joint measurement of the whole system is sometimes necessary. Thus a fundamental question in quantum information is when does optimal extraction of information from a composite system demands access to the whole and when does it not.
@@
This question frequently arises in many quantum information theoretic protocols of practical importance, such as quantum change-point detection [8, 9, 10], multiple-copy state discrimination [11], and quantum key distribution [12]. In all these scenarios, essentially, the objective is to determine the state of an unknown quantum sequence $\left(\rho_{1},\dots,\rho_{k}\right)$ whose members $\rho_{i}$ are drawn from ensembles of which we have complete knowledge. For a given sequence of finite length, this boils down to the problem of sequence discrimination, a natural generalization of the well-studied quantum state discrimination problem [13, 14, 15, 16], where one aims to discriminate between the possible states a quantum system could be in. Since the state of a sequence $\left(\rho_{1},\dots,\rho_{k}\right)$ is of the form $\rho_{1}\otimes\cdots\otimes\rho_{k}$, the question is whether the optimal measurement, i.e., optimal according to some well-defined figure of merit, is local, i.e., comprising measurements on the individual members, or collective, i.e., requiring joint measurement(s). Moreover, even when local access is sufficient, it is necessary to know whether coordinated or adaptive strategies provide any advantage over fixed strategies that do not involve adaptation based on the outcomes of already performed measurements. In fact, we know of instances where the optimal measurement is local and fixed [17], local and adaptive [18], and collective [19]; in particular, in the quantum change-point problem, it is collective [8], whereas in multiple-copy discrimination, depending on the problem specification, it could be either of the two [18, 19].
@@
Unfortunately, given a sequence discrimination problem, we do not have any definite characterization or prescription that could, at the very least, tell us about the nature of the optimal measurement. As it is, discriminating between states of a composite quantum system is known to be challenging [5, 7, 6] and for sequences more so as they could be of any finite length and also come in different varieties: they could be independent and identically distributed (i.i.d.), where the members of a sequence are independent (i.e., selection of an individual member does not depend on what states have already appeared in the earlier positions) and identically distributed (i.e., the members are drawn from the same ensemble), or non-i.i.d., where members are either drawn independently but from different ensembles or not independently at all.
@@
In this paper, we take a significant step towards addressing whether an optimal measurement is local or collective only from the problem description. We prove that if the members of a sequence are secretly and independently selected either from the same ensemble or even from different ensembles, the optimal measurement is local; in particular, it constitutes fixed measurements on the individual elements. The result holds for both minimum-error and unambiguous state discrimination paradigms. For the latter, however, additional assumptions are required for a nontrivial result since unambiguous discrimination applies to states that satisfy certain conditions [20, 21].
@@
Note that as long as the members of a given sequence are selected independently, irrespective of the probability distribution associated with the selection being identical or different, the optimal measurement to determine its state is always local and fixed. Therefore, if the independence condition does not hold, all types of measurements mentioned earlier, are possible. Future works, therefore, only need to address the optimality question in the non-independent scenario.
II Problem statement and main result
@@
Suppose we are given an unknown sequence of length $k\in\mathbb{N}$ and wish to determine its state as well as possible. The members of the given sequence are secretly and independently drawn from potentially different but known ensembles. The ensemble from which the $i$th member is drawn is labeled as $\mathcal{E}^{i}$, where
@@
with $\rho_{j}^{i}$ being density operators on $\mathbb{C}^{d_{i}}$ for $d_{i}\geqslant 2$. Thus, for a sequence of length $k$, we have $k$ such ensembles $\mathcal{E}^{1},\dots,\mathcal{E}^{k}$. In this way, we can account for all possibilities: (i) every member is drawn from a different ensemble; some are drawn from the same ensemble, in which case, though the labeling of ensembles is different, they are, in fact, identical; and (iii) every member is drawn from the same ensemble, so here no indexing of ensembles would be necessary. Note that the last one corresponds to i.i.d. sequences, whereas every other case corresponds to independent but not identically distributed sequences.
Thus, given a sequence of length $k$, we only know that the $i$th member is drawn from $\mathcal{E}^{i}$ and $\eta_{j}^{i}$ is the prior probability of it being $\rho_{j}^{i}$. With this, the ensemble of all possible sequences of length $k$ is given by
@@
The optimality of a measurement subject to a given set of states depends on the choice of measurement strategy. Here, we consider both minimum-error and unambiguous state discrimination paradigms. The former minimizes the average error and applies to any set of states. The corresponding measure is the success probability, the maximum probability that the unknown state is correctly determined [22, 23]. The latter [24, 25, 26], however, applies to sets of states that satisfy a specific condition, and if this condition is met, this approach correctly determines the unknown state with a nonzero probability; for example, a set of pure states can be unambiguously discriminated if and only if they are linearly independent [20]. Closed-form solutions for both are known in the two-state case [23, 27] and in specific instances where the states satisfy certain symmetry properties [28]. For generic ensembles, finding solutions is generally hard; however, semidefinite programs exist [29, 30], and the optimum success probability can be obtained as the output of such programs.
The main result of this paper is the following theorem.
@@
Let $p\left(\mathcal{E}\right)$ denote the optimum probability for minimum-error or unambiguous discrimination between the elements of an ensemble $\mathcal{E}$. Then,
$p\left(\mathcal{E}_{k}\right)=\prod_{i=1}^{k}p\left(\mathcal{E}^{i}\right).$ (3)
Thus the optimal measurement for discriminating between the elements of $\mathcal{E}_{k}$ is local and comprises of individual measurements on the members of the sequence. In particular, for the $i$th member, the local measurement corresponds to the optimal measurement that discriminates between the elements of the ensemble $\mathcal{E}^{i}$.
Since minimum-error discrimination is possible for any set of states, one would always have a nonzero success probability. Specifically for sequences, this means that for any set of ensembles $\{\mathcal{E}^{i}\}$, we have $p(\mathcal{E}^{i})>0$ for all $i$ and as a result $p(\mathcal{E}_{k})>0$ in Eq. (3). However, for unambiguous discrimination this will not be the case in general. Therefore, our result shows that for unambiguous discrimination of a set of quantum sequences, a nontrivial result, i.e., $p\left(\mathcal{E}_{k}\right)>0$, is obtained if and only if for each ensemble $\mathcal{E}^{i},p(\mathcal{E}^{i})>0$, that is, the corresponding ensembles can be unambiguously discriminated.
It is useful to compare our main result with a couple of previous ones obtained under additional assumptions. The authors in [31] studied minimum-cost measurements for quantum states, which is a generalization of minimum-error discrimination. The average cost of the measurement $M=\{M_{i}\}_{i=1}^{N}$ for the ensemble of states $\{(q_{i},\sigma_{i})\}_{i},\;i\in[N]$ with respect to the cost matrix $C=[C_{ij}]$ is denoted $\tilde{C}(M)$ and defined as
$\tilde{C}(M)=\sum_{ij}q_{i}C_{ij}\operatorname{Tr}(\sigma_{i}M_{j}).$ (4)
The minimum cost is computed by taking the optimal measurement that minimizes this average
$\tilde{C}_{\mathrm{min}}=\min_{\{M\}}\tilde{C}(M).$ (5)
Minimum-error discrimination corresponds to the case where $C_{ij}=1-\delta_{ij}.$ For a sequence of quantum states, there is a local minimum-cost problem for each component state as well as a global minimum-cost problem associated with the sequence as a whole. The authors in [31] considered sequences of i.i.d. quantum states where the local cost matrices are identical for the different component states. They proved that under the assumption that the global cost matrix is a linear function of the local cost matrices, the minimum-cost measurement for the sequence is given by the tensor product of the local minimum-cost measurements. In this paper, in contrast, we show that neither the assumption of identical distribution for the component states nor an assumption on the relationship between local and global cost matrices is necessary for proving the optimality of local measurements; the mere independence of the component states is sufficient to guarantee that.
The optimality of fixed local measurements for unambiguous discrimination of quantum sequences were proved in [17] with a number of additional assumptions. There, the states of the sequence were drawn with equal probability from a set of linearly independent pure quantum states whose pairwise inner products were real and equal. Thus the component states of the sequences were identically distributed and also satisfied highly symmetric conditions due to the constraints on their prior probabilities and mutual overlaps. In this paper, we show that one does not need these additional assumptions for the optimality of local measurements; their independence is sufficient to guarantee that.
In the following sections, we prove Theorem 1 for the minimum-error and unambiguous paradigm.
III Minimum-error discrimination of quantum sequences
We need the following two results to prove Theorem 1 for this case. The first one is the Holevo–Yuen–Kennedy–Lax Theorem [32, 33], which provides a necessary and sufficient condition for the optimality of a minimum-error discrimination measurement for a given ensemble (note that the optimal measurement is not unique [34]). We only state the theorem here; the proof can be found in [33].
Theorem 2 (Holevo–Yuen–Kennedy–Lax [32, 33]).
@@ III Minimum-error discrimination of quantum sequences
The second result we need is the following lemma.
@@
Let $A,B,C,$ and $D$ be positive semidefinite operators such that $A\succeq C$ and $B\succeq D$. Then, $A\otimes B\succeq C\otimes D$.
Proof.
$=(A\otimes(B-D))+((A-C)\otimes D).$
Since $A-C$ and $B-D$ are positive semidefinite, their tensor products with any positive semidefinite operator also yield positive semidefinite operators. Hence, $A\otimes(B-D)$ and $(A-C)\otimes D$ are positive semidefinite. The sum of two positive semidefinite operators remains positive semidefinite. Therefore, $A\otimes B-C\otimes D$ is positive semidefinite, and thus $A\otimes B\succeq C\otimes D$. ∎
@@
Without loss of generality, assume that an optimal measurement for minimum-error discrimination between the elements of the ensemble $\mathcal{E}^{i}$ is $M^{i}=\{M_{x}^{i}\}_{x=1}^{\ell_{i}},\text{ where }M^{i}_{x}\in\operatorname{Pos}(\mathbb{C}^{d_{i}})\text{ and }\sum_{x=1}^{\ell_{i}}M_{x}^{i}=\mathbb{1}_{d_{i}}$ for all $x\in[\ell_{i}]$ and $i\in[k]$, where $\operatorname{Pos}(\mathbb{C}^{d_{i}})$ denotes the set of positive semidefinite operators on $\mathbb{C}^{d_{i}}$. It follows that
@@
$=$ $\mathbb{1}_{d_{1}}\otimes\cdots\otimes\mathbb{1}_{d_{k}}$
showing that $\{M^{1}_{x_{1}}\otimes\cdots\otimes M^{k}_{x_{k}}\}$ is a bona fide measurement on $\mathcal{E}_{k}$. This, together with Eq. (9), demonstrates that the measurement whose elements are ${M^{1}_{x_{1}}\otimes\cdots\otimes M^{k}_{x_{k}}}$ is indeed an optimal measurement for discriminating the elements of the ensemble $\mathcal{E}_{k}$, as established by Theorem 2.
@@
IV Unambiguous discrimination of quantum sequences
Unlike the minimum-error case, for nontrivial unambiguous discrimination, a set of quantum states has to satisfy a certain condition. In Sec. IV.1, we derive the condition under which a set of quantum sequences can be unambiguously discriminated. We then present the proof of Theorem 1 in Sec. IV.2, which gives an explicit formula of $p(\mathcal{E}_{k})$ in terms of $\{p(\mathcal{E}^{i})\}$.
IV.1 Conditions for unambiguous discrimination of quantum sequences
Let $Q=\{\sigma_{1},\ldots,\sigma_{N}\}$ be a set of quantum states and $\text{supp}(Q)$ denote the Hilbert space spanned by the eigenvectors of the matrices $\{\sigma_{1},\ldots,\sigma_{N}\}$ that correspond to nonzero eigenvalues. Additionally, let $S(\mathcal{E})$ represent the set of states of an ensemble $\mathcal{E}$.
@@ IV.1 Conditions for unambiguous discrimination of quantum sequences
(from Ref. [21]) The set of density operators $Q=\{\sigma_{1},\ldots,\sigma_{N}\}$ can be unambiguously discriminated if and only if $\text{supp}(Q)\neq\text{supp}(Q_{i})$, where $Q_{i}=Q\setminus\{\sigma_{i}\}$ for all $i\in[N]$. If $\sigma_{i}$ are rank-$1$ operators (pure states), then this condition is same as the set $Q$ being linearly independent.
Since we now want to discriminate the set of sequences unambiguously, they must satisfy Lemma 4. That is, for $S(\mathcal{E}_{k})$ to be unambiguously discriminable, it must hold that
@@
for all $x_{i}\in[\ell_{i}]$ and $i\in[k]$. In the following lemma, we show that a set of quantum sequences satisfies this condition if and only if the individual ensembles do.
@@
We will need the following two lemmas to prove Lemma 5.